# DISTA-Net: Dynamic Closely-Spaced Infrared Small Target Unmixing

### v1

@@ 4.2 DISTA-Net: A Dynamic Framework
To integrate ISTA with network-based sparse representation, DISTA-Net maps each ISTA iteration to a network layer. By stacking these layers and enabling end-to-end training, the network optimizes parameters automatically, improving upon traditional ISTA’s manually designed parameters while maintaining interpretability. Building upon ISTA-Net’s framework for compressed sensing, we introduce two key improvements. First, we design a data-adaptive nonlinear transformation function $\mathcal{F}_{d}(\cdot)$ that maps images to richer dimensional representations while emphasizing significant image regions. The transformation follows:
@@
where $\operatorname{Soft}(\cdot,\theta_{d})$ denotes soft-thresholding with learnable parameter $\theta_{d}$, and $k$ is the stage index. The left inverse $\tilde{\mathcal{F}}$ satisfies ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with a structure symmetric to $\mathcal{F}_{d}(\cdot)$. This yields:
@@
Second, we introduce a dynamic threshold module that adapts to input image variations, addressing the sensitivity of sparse vector perturbations in image generation. This flexible thresholding mechanism improves upon fixed parameters that can be either too strict or too lenient. As shown in Fig. 4, DISTA-Net comprises $N$ stages. Each stage $k$ ($k>1$) contains three components: $\mathcal{F}_{d}^{(k)}$, $\operatorname{Soft}(\cdot,\theta_{d}^{(k)})$, and ~$\tilde{\mathcal{F}}^{(k)}$. The input $\mathbf{r}^{(k)}$ is derived from $\tilde{\mathbf{s}}^{(k-1)}$ via Eq. (7), processed through these components sequentially to generate $\tilde{\mathbf{s}}^{(k)}$ for the next iteration.
Dynamic Transform.
While trainable non-linear transformations overcome limitations of handcrafted methods, their fixed post-training parameters result in static transformation patterns. We address this by introducing a dual-branch Dynamic Transform module $\mathcal{F}_{d}^{(k)}$ at the $k$-th stage (Fig. 4).
@@
$W=f(\tilde{\mathbf{s}}^{(k-1)})$ (13)
The Dynamic Conv module(Fig. 4) applies this weight vector as an adaptive convolutional kernel to $\mathbf{r}^{(k)}$:
$w_{r}=C(W,\mathbf{r}^{(k)})$ (14)
@@
Dynamic Soft-Thresholding. Unlike ISTA-Net’s fixed threshold $\theta$, we propose a Dynamic Soft-Thresholding module that adapts $\theta_{d}$ based on input image information. This approach better handles densely overlapped targets and spatial context variations.
As shown in Fig. 4, the module employs dual convolutional layers to capture multi-scale features as in Li et al. (2024a, 2019a). The $\mathcal{F}_{d}^{(k)}$ output passes through two $3\times 3$ convolutions, generating feature maps $\tilde{U}_{1}$ and $\tilde{U}_{2}$. These are concatenated to form $\tilde{U}=[\tilde{U}_{1},\tilde{U}_{2}]$. Spatial relationships are captured through parallel pooling operations:
$SA_{\text{avg}}=P_{\text{avg}}(\tilde{U}),\quad SA_{\text{max}}=P_{\text{max}}(\tilde{U}).$ (16)
@@
The dynamic threshold $\theta_{d}$ is then computed by combining these masks with multi-scale feature maps through a final convolution.
@@
Here, $\mathcal{L}_{\text{discrepancy}}$ computes the MSE between reconstructed $\tilde{s}_{i}^{(N)}$ and ground truth $s_{i}$, while $\mathcal{L}_{\text{constraint}}$ enforces multi-stage identity constraints through ~$\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, with $\gamma=0.01$ balancing these objectives.
@@ 5 Experiments
ACTNet Zhang et al. (2023) 46.212M 62.80G 45.61 0.38 7.46 41.13 83.12 95.95 35.4526 99.70
CTNet Lim et al. (2017) 0.400M 2.756G 45.11 0.38 7.53 40.39 82.11 95.14 35.1499 99.70
DCTLSA Zeng et al. (2023) 0.865M 13.69G 44.51 0.39 7.35 39.35 81.15 94.34 34.6314 99.65
EDSR Lim et al. (2017) 1.552M 12.04G 45.32 0.33 7.07 40.58 83.24 95.41 35.3724 99.71
EGASR Qiu et al. (2023) 2.897M 17.73G 45.51 0.42 8.03 41.32 85.71 95.08 34.5681 99.66
FeNet Wang et al. (2022) 0.348M 2.578G 45.77 0.42 8.19 42.13 83.30 94.80 34.1531 99.66
RCAN Zhang et al. (2018a) 1.079M 8.243G 45.87 0.42 7.96 41.81 83.61 95.57 35.2119 99.69
RDN Zhang et al. (2018b) 22.306M 173.0G 45.81 0.35 7.11 41.07 84.07 96.43 36.4686 99.74
SAN Dai et al. (2019) 4.442M 34.05G 45.95 0.36 7.35 41.17 84.32 96.57 36.5037 99.74
SRCNN Dong et al. (2015) 0.019M 1.345G 29.06 0.23 4.10 21.65 49.95 69.39 28.7608 98.44
SRFBN Li et al. (2019b) 0.373M 3.217G 46.05 0.43 8.31 42.83 83.72 94.95 34.0174 99.68
HAN Niu et al. (2020) 64.342M 495.0G 45.70 0.39 7.46 40.90 83.61 96.17 35.2703 99.71
HiT-SIR Zhang et al. (2024) 0.769M 7.357G 44.51 0.40 7.16 39.55 80.94 94.50 34.7400 99.70
HiT-SNG Zhang et al. (2024) 0.952M 13.324G 45.01 0.39 7.34 40.19 81.98 95.17 35.1390 99.71
HiT-SRF Zhang et al. (2024) 0.844M 7.945G 44.71 0.40 7.71 40.03 80.94 94.49 34.5449 99.67
@@
ISTA-Net Zhang and Ghanem (2018) 0.171M 12.77G 45.16 0.41 7.71 40.57 82.58 94.53 33.9215 99.68
ISTA-Net+ Zhang and Ghanem (2018) 0.337M 24.33G 46.06 0.42 7.66 41.58 84.46 96.17 36.0892 99.72
LAMP Daubechies et al. (2004) 2.126M 0.278G 14.22 0.05 1.11 7.31 21.56 41.06 27.8299 96.89
LIHT Daubechies et al. (2004) 21.10M 1.358G 10.35 0.06 0.92 4.99 14.74 30.5 27.5107 96.42
LISTA Daubechies et al. (2004) 21.10M 1.358G 30.13 0.25 4.13 22.29 51.18 72.82 29.8936 99.12
FISTA-Net Xiang et al. (2021) 0.074M 18.96G 44.66 0.45 7.68 39.74 81.24 94.19 35.7519 99.67
TiLISTA Daubechies et al. (2004) 2.126M 0.278G 14.95 0.06 1.23 7.72 22.50 46.23 27.7038 97.40
$\star$ DISTA-Net (Ours) 2.179M 35.10G 46.74 0.38 7.54 42.44 86.18 97.14 37.8747 99.79
@@ 5.1 Experimental Settings
Training. Using images from the CSIST dataset as input, we employ the method mentioned in  4.1 to perform sub-pixel division with a sampling grid ratio of $c$, generating an unmixed high-resolution grid as ground truth. Specifically, for the $i$-th target with parameters $(x_{i},y_{i},g_{i})$, the intensity $g_{i}$ is assigned to the pixel at position $\left(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2}\right)$, with other pixels set to zero. With an appropriate selection of c, each target is resolved as a distinct point in the unmixed grid, enabling precise spatial separation.
Testing. We project the unmixed high-resolution grid back to the original 11×11 coordinate space to calculate the CSO-mAP against the ground truth target positions. Post-processing is applied to the model’s predicted images, where pixels with intensity values above a threshold of 50 are retained as predicted targets. For the i-th predicted target, its coordinates are computed as $\left(\frac{x_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c},\frac{y_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c}\right)$. To provide a more comprehensive assessment of the model’s performance, we also compute PSNR and SSIM metrics by directly comparing the predicted high-resolution images with the sub-pixel divided ground truth images.
Hyperparameters. We set $c=3$ as the baseline. The network is trained on randomly selected batches of 64 images, with the DIST-Net configured with $N=6$ stages and the Dynamic Transform branch coefficient $(1-\alpha)$ set to 0.3.
5.2 Comparison with SOTA Methods
@@
For localization accuracy, we adopt CSO-mAP with different distance thresholds (from AP-05 to AP-25), where a smaller threshold indicates stricter localization precision requirement. For instance, AP-05 evaluates the detection accuracy within 0.05-pixel distance, representing an extremely high precision demand. Under practical scenarios where 0.25-pixel precision (AP-25) is sufficient, DISTA-Net achieves 97.14% accuracy, outperforming most existing methods. Even under more stringent conditions (AP-15 and AP-10), our method maintains competitive performance with 42.44% and 7.54% accuracy respectively, demonstrating its robust localization capability across different precision requirements.
In terms of model efficiency, DISTA-Net achieves these results with moderate computational costs (2.179M parameters and 35.103G FLOPs), showing better efficiency compared to methods like ACTNet (46.212M, 62.798G) and HAN (64.342M, 0.495T). Additionally, DISTA-Net demonstrates superior image quality with the highest PSNR (37.8747) and SSIM (99.79) scores among all compared methods, indicating its excellent capability in preserving image details while performing target localization.
@@
Figure 5: Visual comparison of 3× sub-pixel division reconstruction for scenes containing different numbers of adjacent targets. The red boxes highlight regions where targets exhibit significant sub-pixel characteristics.
Visual comparison. Fig. 5 compares reconstruction results with 3× sub-pixel division across methods for scenes containing 3–5 adjacent targets, assessing performance in dense multi-target scenarios. Existing methods struggle to reconstruct closely-spaced targets, exhibiting blurred boundaries or merged detections, with degradation worsening for higher target counts (e.g., five-target cases). In contrast, DISTA-Net preserves both target quantity and sub-pixel positions while maintaining sharp boundaries and accurate spatial distributions, even under extreme density.
@@ 5.3 Ablation Study
Method #P $\downarrow$ FLOPs $\downarrow$ CSO-mAP
  ISTA-Net 0.171M 12.766G 45.16
DISTA-Net w/o DT 0.171M 39.544G 46.32
DISTA-Net w/o Thres. 5.111M 92.62G 46.17
$\star$ DISTA-Net (Ours) 2.179M 35.103G 46.74
@@
Effect of different components.
We conduct ablation studies to evaluate the contribution of each component in DISTA-Net, with results shown in Table 2. The second row (DISTA-Net w/o DT) corresponds to the model variant without Dynamic Transform and the third row (DISTA-Net w/o Thres.) represents the model without Dynamic Soft-Thresholding. Compared to the baseline ISTA-Net, our complete model shows notable improvements in both CSO-mAP (45.16% to 46.74%) and AP-20 (82.58% to 86.18%).
@@
ISTA-Net Zhang and Ghanem (2018) 0.171M 39.544G 66.90
ISTA-Net+ Zhang and Ghanem (2018) 0.225M 48.158G 68.50
CFGN Dai et al. (2023a) 0.538M 4.122G 67.95
$\star$ DISTA-Net (Ours) 5.153M 102.4G 69.58
@@
ISTA-Net Zhang and Ghanem (2018) 0.171M 89.51G 71.19
ISTA-Net+ Zhang and Ghanem (2018) 0.225M 103.0G 71.09
CFGN Dai et al. (2023a) 0.548M 4.202G 70.38
$\star$ DISTA-Net (Ours) 6.409M 142.3G 72.84
@@
Model Performance and Sampling Grid. To investigate the impact of sampling grid ratio on model performance and computational efficiency, we conduct experiments with other different sampling grid ratios (c = 5, and 7) across various methods. Table 3 presents a comprehensive comparison of model performance under different sampling configurations.
As the sampling grid ratio increases, we observe a general trend of improved detection performance across all methods, which is attributed to the higher precision in target position representation. DISTA-Net consistently maintains a leading advantage in positioning accuracy, outperforming other methods at all configurations. However, this performance gain comes at the cost of substantially increased computational complexity. Notably, DISTA-Net demonstrates better efficiency in utilizing the increased resolution. Compared to ISTA-Net and ISTA-Net+, which show similar trends but lower performance, our method achieves higher accuracy gains with the same increase in sampling ratio. Based on these observations, we suggest selecting an appropriate sampling grid ratio according to real-world requirements to achieve a good balance between detection accuracy and computational efficiency.
Hyperparameters Analysis. We analyzed model performance versus stage numbers and dynamic branch coefficients in Supplementary, demonstrating robust architectural and parametric design.

### v2

@@ 4.2 DISTA-Net: A Dynamic Framework
Algorithm 1 DISTA-Net
0: CSIST image $\mathbf{z}$, steering matrix $\mathbf{G}(\mathbf{x},\mathbf{y})$, initial matrix $Q_{\text{init}}$, number of stages NN, step size $\{\rho^{(k)}\}_{k=1}^{N}$
[…]
2:for $k=1$ to NN do
3:  $\mathbf{r}^{(k)}\leftarrow\tilde{\mathbf{s}}^{(k-1)}-\rho^{(k)}\mathbf{G}^{\top}(\mathbf{G}\tilde{\mathbf{s}}^{(k-1)}-\mathbf{z})$
4:  $\acute{\mathbf{s}}^{(k)}\leftarrow\mathcal{F}_{d}(\tilde{\mathbf{s}}^{(k-1)},\mathbf{r}^{(k)})$
5:  $\theta_{d}^{(k)}\leftarrow\mathrm{DTG}^{(k)}(\acute{\mathbf{s}}^{(k)})$
6:  $\tilde{\mathbf{s}}^{(k)}\leftarrow\tilde{\mathcal{F}}(\operatorname{Soft}(\mathcal{F}_{d}(\mathbf{r}^{(k)}),\theta_{d}^{(k)}))$
7:end for
The details of our DISTA-Net are illustrated in Fig.  4. Building upon ISTA-Net’s framework, we introduce two key improvements. First, we design a data-adaptive nonlinear transformation function $\mathcal{F}_{d}(\cdot)$ that maps images to richer dimensional representations while emphasizing significant image regions. The transformation follows:
@@
where $\operatorname{Soft}(\cdot,\theta_{d})$ denotes soft-thresholding with learnable parameter $\theta_{d}$, and $k$ is the stage index. The left inverse $\tilde{\mathcal{F}}$ satisfies $\tilde{\mathcal{F}}(\cdot)\circ\mathcal{F}_{d}(\cdot)=\mathbf{I}$, without requiring structural symmetry to $\mathcal{F}_{d}(\cdot)$, yielding:
@@
Second, we introduce a dynamic threshold module that adapts to input image variations, addressing the sensitivity of sparse vector perturbations in image generation. This flexible thresholding mechanism improves upon fixed parameters that can be either too strict or too lenient. As shown in Fig. 4, DISTA-Net comprises $N$ stages. Each stage $k$ ($k>1$) contains three components: $\mathcal{F}_{d}^{(k)}$, $\operatorname{Soft}(\cdot,\theta_{d}^{(k)})$, and $\tilde{\mathcal{F}}^{(k)}$. The input $\mathbf{r}^{(k)}$ is derived from $\tilde{\mathbf{s}}^{(k-1)}$ via Eq. (7), processed through these components sequentially to generate $\tilde{\mathbf{s}}^{(k)}$ for the next iteration. Overall, the proposed DISTA-Net can be referenced in Algorithm  1.
Dynamic Transform. While trainable non-linear transformations overcome limitations of handcrafted methods, their fixed post-training parameters result in static transformation patterns. We address this by introducing a dual-branch Dynamic Transform module $\mathcal{F}_{d}^{(k)}$ at the $k$-th stage.
@@
$W=f(\tilde{\mathbf{s}}^{(k-1)}).$ (13)
The Dynamic Conv module applies this weight vector as an adaptive convolutional kernel to $\mathbf{r}^{(k)}$:
$w_{r}=C(W,\mathbf{r}^{(k)}).$ (14)
@@
Dynamic Soft-Thresholding. Unlike ISTA-Net’s fixed threshold $\theta$, we propose a Dynamic Thresholding Generator(DTG) module that adapts $\theta_{d}$ based on input image information. This approach better handles densely overlapped targets and spatial context variations.
As shown in Fig. 4, the module employs dual convolutional layers to capture multi-scale features as in [16, 12]. The $\mathcal{F}_{d}^{(k)}$ output passes through two $3\times 3$ convolutions, generating feature maps $\tilde{U}_{1}$ and $\tilde{U}_{2}$. These are concatenated to form $\tilde{U}=[\tilde{U}_{1},\tilde{U}_{2}]$. Spatial relationships are captured through parallel pooling operations:
[…]
@@
The dynamic threshold $\theta_{d}$ is then computed by combining these masks with multi-scale feature maps through a final convolution $C(\cdot)$:
$\theta_{d}=C(\sum_{i=1}^{N}{(\tilde{SA})_{i}\cdot\tilde{U}_{i}}).$ (19)
@@
Here, $\mathcal{L}_{\text{discrepancy}}$ computes the MSE between reconstructed $\tilde{s}_{i}^{(N)}$ and ground truth $s_{i}$, while $\mathcal{L}_{\text{constraint}}$ enforces multi-stage identity constraints through $\tilde{\mathcal{F}}^{(k)}\circ\mathcal{F}_{d}^{(k)}\approx\mathbf{I}$, with $\gamma=0.01$ balancing these objectives.
@@ 5 Experiments
ACTNet [45] 46.212M 62.80G 45.61 0.38 7.46 41.13 83.12 95.95 36.8879 0.9854
CTNet [30] 0.400M 2.756G 45.11 0.38 7.53 40.39 82.11 95.14 35.6154 0.9833
DCTLSA [38] 0.865M 13.69G 44.51 0.39 7.35 39.35 81.15 94.34 35.2259 0.9832
EDSR [19] 1.552M 12.04G 45.32 0.33 7.07 40.58 83.24 95.41 36.4349 0.9816
EGASR [25] 2.897M 17.73G 45.51 0.42 8.03 41.32 85.71 95.08 35.4549 0.9745
FeNet [31] 0.348M 2.578G 45.77 0.42 8.19 42.13 83.30 94.80 36.4507 0.9864
RCAN [43] 1.079M 8.243G 45.87 0.42 7.96 41.81 83.61 95.57 36.5860 0.9831
RDN [44] 22.306M 173.0G 45.81 0.35 7.11 41.07 84.07 96.43 36.1883 0.9840
SAN [2] 4.442M 34.05G 45.95 0.36 7.35 41.17 84.32 96.57 37.1805 0.9848
SRCNN [8] 0.019M 1.345G 29.06 0.23 4.10 21.65 49.95 69.39 29.4140 0.9566
SRFBN [17] 0.373M 3.217G 46.05 0.43 8.31 42.83 83.72 94.95 34.2945 0.9815
HAN [23] 64.342M 495.0G 45.70 0.39 7.46 40.90 83.61 96.17 36.6048 0.9826
HiT-SNG [42] 0.952M 13.324G 45.01 0.39 7.34 40.19 81.98 95.17 35.4038 0.9805
@@
ISTA-Net [40] 0.171M 12.77G 45.16 0.41 7.71 40.57 82.58 94.53 35.6733 0.9862
ISTA-Net+ [40] 0.337M 24.33G 46.06 0.42 7.66 41.58 84.46 96.17 38.4998 0.9887
LAMP [7] 2.126M 0.278G 14.22 0.05 1.11 7.31 21.56 41.06 27.8181 0.9405
LIHT [7] 21.10M 1.358G 10.35 0.06 0.92 4.99 14.74 30.50 27.5657 0.9198
LISTA [7] 21.10M 1.358G 30.13 0.25 4.13 22.29 51.18 72.82 30.1285 0.9537
FISTA-Net [34] 0.074M 18.96G 44.66 0.45 7.68 39.74 81.24 94.19 38.4773 0.9872
TiLISTA [7] 2.126M 0.278G 14.95 0.06 1.23 7.72 22.50 46.23 27.7790 0.8895
$\star$ DISTA-Net (Ours) 2.179M 35.10G 46.74 0.38 7.54 42.44 86.18 97.14 38.3825 0.9887
@@ 5.1 Experimental Settings
Training. Using CSIST-100K images as input, we apply Sec. 4.1’s method to perform sub-pixel division with a sampling grid ratio of $c$. This generates an unmixed high-resolution grid as ground truth, where for each target $(x_{i},y_{i},g_{i})$, the intensity $g_{i}$ is assigned to the pixel at position $\left(c\cdot x_{i}+\frac{c-1}{2},c\cdot y_{i}+\frac{c-1}{2}\right)$ while other pixels remain zero. The selected $c$ value ensures each target appears as a distinct point, enabling accurate spatial separation.
Testing. We (1) apply post-processing with an intensity threshold of 50 to identify predicted targets; (2) project the unmixed grid back to the original $11\times 11$ space via $\left(\frac{x_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c},\frac{y_{i}-\left\lfloor\frac{c-1}{2}\right\rfloor}{c}\right)$. These mapped coordinates are then compared with ground truth target positions to compute CSO-mAP. Additionally, we calculate PSNR and SSIM by directly comparing the full predicted and ground truth high-resolution images.
Hyperparameters. Our configuration uses: $c=3$ (baseline grid ratio), batch size of 64, DISTA-Net with 6 stages ($N=6$), and Dynamic Transform branch coefficient $(1-\alpha)=0.3$.
5.2 Comparison with State-of-the-Art Methods
@@
For localization accuracy, we adopt CSO-mAP with different distance thresholds (from AP-05 to AP-25), where a smaller threshold indicates a stricter localization precision requirement. For instance, AP-05 evaluates the detection accuracy within 0.05-pixel distance, representing an extremely high precision demand. Considering the inherent error of 0.236 pixel width at subpixel division factor $c=3$, AP-20 and AP-25 (with 0.20 and 0.25 precision requirements respectively) serve as primary performance benchmarks. DISTA-Net achieves remarkable accuracy rates of 86.18% and 97.14% accuracy, outperforming most existing methods. On the mAP metric reflecting average CSIST unmixing performance, our method maintains a leading advantage with 46.74% accuracy, demonstrating its robust localization capability across different precision requirements.
In terms of model efficiency, DISTA-Net achieves these results with moderate computational costs (2.179M parameters and 35.103G FLOPs), showing better efficiency compared to methods like ACTNet (46.212M, 62.798G) and HAN (64.342M, 0.495T). Additionally, DISTA-Net achieves a leading SSIM score of 0.9887 and a satisfactory PSNR of 38.3825, effectively preserving image details.
@@
Visual comparison. Fig. 5 compares reconstruction results with $3\times$ sub-pixel division across methods for scenes containing $3\sim 5$ targets, assessing performance in dense multi-target scenarios. Existing methods struggle to reconstruct closely spaced targets, exhibiting blurred boundaries or merged detections, with degradation worsening for higher target counts (e.g., five-target cases). In contrast, DISTA-Net preserves both target quantity and sub-pixel positions while maintaining sharp boundaries and accurate spatial distributions, even under extreme density.
Figure 5: Visual comparison of $3\times$ sub-pixel division reconstruction for scenes containing different numbers of closely-spaced infrared small targets. The red boxes highlight regions where targets exhibit significant sub-pixel characteristics.
@@ 5.3 Ablation Study
Method CSO-mAP AP-05 AP-10 AP-15 AP-20 AP-25
  ISTA-Net [40] 45.16 0.41 7.71 40.57 82.58 94.53
DISTA-Net w/o DT 46.32 0.34 6.83 40.76 86.18 97.50
DISTA-Net w/o Thres. 46.17 0.44 7.77 42.18 84.67 95.79
$\star$ DISTA-Net (Ours) 46.74 0.38 7.54 42.44 86.18 97.14
@@
Effect of different components. We conduct ablation studies to evaluate the contribution of each component in DISTA-Net, with results shown in Table 2. The second row (DISTA-Net w/o DT) corresponds to the model variant without Dynamic Transform and the third row (DISTA-Net w/o Thres.) represents the model without Dynamic Soft-Thresholding. Compared to the baseline ISTA-Net, our complete model shows notable improvements in both CSO-mAP (45.16% to 46.74%) and AP-20 (82.58% to 86.18%).
@@
ISTA-Net  [40] 0.171M 39.544G 66.90 56.73 87.26
ISTA-Net+  [40] 0.225M 48.158G 68.50 57.96 89.52
CFGN  [3] 0.538M 4.122G 67.95 58.08 88.35
$\star$ DISTA-Net (Ours) 5.153M 102.4G 69.58 60.95 90.85
@@
ISTA-Net  [40] 0.171M 89.51G 71.19 76.45 84.16
ISTA-Net+  [40] 0.225M 103.0G 71.09 74.90 84.90
CFGN  [3] 0.548M 4.202G 70.38 73.88 83.97
$\star$ DISTA-Net (Ours) 6.409M 142.3G 72.84 78.47 86.09
@@
Model Performance and Sampling Grid. We evaluate sampling ratios $c=5$ and $7$ across methods, analyzing their impact on performance and efficiency. As the sampling grid ratio increases, all methods show improved detection performance due to better target positioning precision. DISTA-Net consistently outperforms other methods across all configurations. While ISTA-Net and ISTA-Net+ exhibit similar trends with lower overall performance, our proposed method achieves greater accuracy improvements for equivalent increases in sampling ratio. This advantage is particularly pronounced in the AP-10 ($c=5$) and AP-15 ($c=7$) metrics (see Table 3 for complete results). However, these performance gains are accompanied by substantially increased computational complexity. We recommend selecting an appropriate sampling grid ratio that aligns with specific application requirements to achieve an optimal balance between detection accuracy and computational efficiency.
Ours vs Super-Resolution + Detector Pipeline. The unmixing stage typically serves as a refinement step following the detection phase. For experimental rigor, we conducted experiments implementing the “SR + Detector” pipeline. We used YOLOv11 as the detector following leading SR methods (retrained on our IR data with unmixing GT for point-source super-resolution). The results demonstrate DISTA-Net’s continued advantage: DISTA-Net + YOLOv11 achieved a CSO-mAP of 47.82, outperforming SRFBN + YOLOv11 (45.74) and CFGN + YOLOv11 (46.71). Visual analysis reveals that both conventional SR methods and our unmixing approach generate high-resolution images with well-resolved peaks, enabling effective target separation through simple thresholding. This demonstrates the dominant role of the unmixing stage in this task.
Hyperparameters Analysis. We analyzed model performance versus stage numbers and dynamic branch coefficients in Supplementary, demonstrating robust design.